OpenStudy (anonymous):

12. A sphere has a radius of 4x + 4. Which polynomial in standard form best describes the total surface area of the sphere? Use the formula S=4 3.14r2 for the surface area of a sphere.

OpenStudy (anonymous):

Missing options?

OpenStudy (hoblos):

S=4*3.14r2 = 12.56(4x+4)^2 = 12.56( 16x^2 +32x + 16) = 200.96x^2 +401.92x + 200.96